OpenStudy (anonymous):

Please help So how would you solve this? 5/x-2x=3 ? Please help

OpenStudy (raden):

like this : 

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

multiply by \(x\) to get \[5-2x^2=3x\] then arrange as \[2x^2+3x-5=0\] and solve the quadratic equation

OpenStudy (anonymous):

\[(x-1)(2x+5)=0\] is the factored form, which tells you \[x=2\] or \[x=-\frac{5}{2}\]

OpenStudy (raden):

there is a little typo, maybe. it should be x=1 for the first x

OpenStudy (anonymous):

thanks to both of you. I wish i can give you both medals here. sadly I can't |: